Solve using the addition principle. Dont forget to perform a check
-3+x=21 My answer her is -7 (I don't know if this is right)

Answer by Edwin McCravy(20086)

Solve using the addition principle. Dont forget to perform a check
-3+x=21 My answer her is -7 (I don't know if this is right)

No -7 is not correct.

Here's how to do it

     -3 + x = 21

                   First we get rid of the -3 on the left by adding
                   +3 to both sides. We write +3 under the -3 and
                   also write +3 under the 21 and draw a line:
 
     -3 + x = 21
     +3       +3 
                   
                   Now we add vertically. The -3 and the +3 cancel,
                   so we don't write anything under them.  Bring
                   the x down. Finally, the 21 and the +3 gives 24:

     -3 + x = 21
     +3       +3 
          x = 24

So the solution is 24.

Now lets check by replacing x in the original by 24:

     -3 + x = 21
    -3 + 24 = 21
         21 = 21

Edwin